# Survey of Natural Language Processing for Education: Taxonomy, Systematic Review, and Future Trends

### v3

@@ 3.1 Question Answering
FIGURE 5: The examples of Question Answering in education domain. The red box denotes the given context and the green box denotes the required prediction for the tasks. The displayed examples are extracted from SCIENCEQA and GSM8K, respectively.
Even though QA has been explored as a popular NLP task for decades. Educational QA requires a system to answer the domain-specific questions, related to natural science, social science, language science and so on [109]. These systems act as assistants to help answer difficult questions such that the students could learn from them and boost their performance. Question answering in different subjects may be presented in different formats. In Figure 5, we display two typical examples of question answering in the education domain. To answer the displayed questions, the systems need domain-specific knowledge like mathematics, and science. Furthermore, the contexts of the questions are presented as text and images. Moreover, the output of the questions can presented as an entity or a math expression. In this section, we will present cutting-edge techniques for Textbook QA and MWP solving.
@@ 3.1.1 Textbook Question Answering
Textbook Question Answering (TQA) is a task that requires a system to comprehensively understand the multi-modal information from the textbook curriculum, spreading across text documents, images, and diagrams. The major challenge of textbook question answering is to comprehend the multi-modal domain-specific contexts as well as the questions, and then identify the key information to the questions.
@@
Datasets. Kembhavi et al. [110] introduced TQA dataset, which aims to evaluate a system that contains multi-modal contexts and diverse topics in the scientific domain. Similar datasets such as AI2D [29], DVQA [31], and VLQA [111] have been constructed to support researches on multi-modal reasoning in the scientific domain. However, these datasets lack of annotated explanations for the answers in the form of supporting facts. SCIENCEQA [1] is a large-scale textbook question answering dataset with annotated lectures and explanations. The dataset is collected from elementary and high school science curricula, which contain diverse science topics including natural science, social science, and language science. A recently released TheoremQA dataset further covers textbook questions at the university level [2]. Besides the scientific domain, there are some datasets focusing on the medical domain. MEDQA [112] and MedMCQA [113] are two medical QA datasets encompassing a broad spectrum of healthcare topics deriving from both real-world scenarios and simulated exams.
Methods. From the technical perspective, textbook question answering is visual question answering in nature [114, 115, 116]. Conventional VQA studies employ RNN to encode the question and CNN to encode the image [18, 19], respectively. Then the multi-modal information is fused together to comprehend the questions. Besides, other methods employing spatial attention [20, 21, 22, 23], compositional strategies [24], and bilinear pooling schemes [25, 26] are proposed to boost the performance of VQA.
While there are significant similarities between VQA and textbook question answering, textbook question answering differs in the domain-specific knowledge required for the accompanying context and the innovative inclusion of diagrams and tables. To bridge the gap, Ram et al. [27] proposed a pre-training schema tailored for question answering. Specifically, by masking recurring span selection and selecting the correct span in the passage, their method boosts the performance for textbook question answering when only a hundred examples are available in the specific domains. Adversarial training framework is also adapted to domain generalization [28], which facilitates the question answering models to learn domain-invariant features. Xu et al. [117] introduced a novel Pre-trained Machine Reader as a retrofit of pre-trained Masked Language Models (MLMs), which resolves the discrepancy between model pre-training and down-stream fine-tuning of specific domain MLMs. To understand the diagrams and tables, graph-based parsing methods are developed to extract the concepts from diagrams [29] by converting a diagram to a diagram parse graph. Optical Character Recognition (OCR) is employed to identify the chart-specific answers from the charts, which is further aligned with the questions [30, 31].
Recently, Large Language Models (LLMs) have shown impressive capabilities in various NLP tasks, researchers start applying LLMs to textbook question answering. Lu et al. [1] first applied GPT-3 to solve the textbook question answering. To apply the full advantage of LLMs to image modality, they use caption models to translate visual information into the language modality and prompt LLMs to perform Chain-of-Thought. This inspires the follow-up study [32], distilling the training signals from the LLMs to a smaller model. To inject general knowledge from the LLMs to the specific domains, Wang et al. [33] augmented queries for LLMs on textbook question answering by retrieving evidences from textbooks, such that the response of the LLMs would likely to be authorized by the textbooks. Liu et al. [34] extended the instruction tuning to a multi-modal field, which results in a multi-modal model that connects a vision encoder and an LLM for general-purpose visual and language understanding. More recent studies [35] develop compositional reasoning with LLMs as a planner that assembles a sequence of tools (e.g., Hugging Face as image captioner, Github as text detector, Web Search as bing search, and Python as program verifier.) to execute and answer the textbook questions. Instead of solving the questions with given contexts, Chen et al. [2] introduced a system, that first retrieves the theories from the open domain and then answers the question based on the retrieved theories.
Dataset Science subject Source level Context #Q
TQA life/earth/physical grade 6-8 image/ 26262626K
science science curricula diagram
Geometry3K geometry grade 6-12 digital libraries image/diagram 3333K
AI2D science grade 1-6 image/diagram 5555K
SCIENCEQA natural/social/ grade 1-12 image/diagram 21212121K
language science science curriculum text
MedQA medicine professional medical text 40404040K
board exam
MedMCQA medicine simulated exams text 200200200200K
TheoremQA math/physics/EE/CS university-level exam text $800800$
Dolphin-18K math grade school text 18181818K
DRAW-1K math algebra text 1111K
Math23K math grade school math text 23232323K
MathQA math grade school math text 37373737K
ASDiv math grade school math text 2222K
GSM8K math grade school math text 8888K
IconQA math K-12 math text/image 107107107107K
TABMWP math grade school math text/table 38383838K
TABLE I: Comparison of textbook question answering and MWP solving datasets. The first section includes datasets of Textbook QA. The second section includes datasets of MWP solving. “#Q” denotes the number of questions.
@@ 3.1.2 Math Word Problem solving
Math Word Problem (MWP) solving is a specific type of question answering, aiming to convert a narrative description to an abstract expression. This task is challenging because there remains a wide semantic gap to parse human-readable words into machine-understandable logic.
@@
Datasets. Dolphin-18181818[118] is an early large-scale MWP solving dataset, which contains over 18,0001800018,00018 , 000 annotated math word problems of elementary mathematics. DRAW-1111[119] is introduced as a testbed to algebra word problem solvers. Math23232323[36] is a Chinese MWP solving dataset. Subsequently, MathQA [120] and ASDiv [121] further extend the scale and diversity of the existing datasets, which makes it particularly suited for usage in training deep learning models to solve MWP problems. Cobbe et al. [122] introduced a more challenging dataset GSM8888K, which is complicated in high linguistic diversity. IconQA [123] extends MWP solving to abstract diagram understanding and comprehensive cognitive reasoning by involving diagrams in problems. Moreover, TABMWP [124] includes tables to enhance the difficulty of the MWPs, which requires the table parsing and understanding capabilities [124]. As we can see, there is a set of MWP datasets that examine various mathematics capabilities. With the development of this task, the evaluated datasets become increasingly complicated in linguistic diversity and multi-modalities.
@@
Methods. The investigation of MWP solving can be traced back to 1960196019601960[125]. Early work leverages manually crafted rules and schemas for pattern matching, where human intervention is heavily relied on. Subsequently, various strategies of feature engineering and statistical learning are proposed to boost the performance. With the development of techniques and the emergence of large-scale datasets, deep learning-based methods have been widely applied to MWP solving. The majority of the methods follow an encoder-decoder framework [36], which first encodes the narratives into dense vector space, then generates the mathematical expression token as the output sequence. To improve the generation of complex solution expressions, some methods are proposed to formulate a mathematical expression as an abstract tree [37, 38, 39, 43]. To improve the understanding to the narratives, some studies focus on improving the encoding of the narratives by re-organizing the textual contexts with structure [40, 41]. Due to the weak supervision caused by the lack of annotated expressions, multiple studies [37, 42] employ reinforcement learning to solve MWPs. Also, some researchers [126] apply knowledge distillation to MWP solving tasks, which learns a smaller model from a large pre-trained generic model.
@@
As the MWP solving task is the desirable testbed for evaluating the reasoning capabilities, a series of studies have been developed to discuss the abilities of LLMs for solving MWPs. Researchers find that prompting the LLMs with effective planning strategies could activate the potential of LLMs, which results in impressive results on MWP solving under the zero/few-shot scenarios. Chain-of-thought is a typical promoting strategy that is meant to decompose a complicated MWP to sub-questions such that LLMs could conduct reasoning step by step [127]. Follow-up methods improve it by designing more interactive prompting or constructing more efficient demonstrations [44, 45, 46, 47, 35]. With the emergence of the idea of leveraging LLMs as an agent [128], researchers [48] delegate the computation to a program interpreter, which can be conducted by the LLMs and executed by a Python interpreter. In this case, the reasoning step can be verified in an iteration manner.
@@ 3.2 Question Construction
3.2.1 Question Generation
The goal of QG is to automatically generate a question from a given sentence or paragraph, with the assistance of answer information when available. The challenge lies in identifying the key statement based on the context with or without the utilization of the answer (namely answer-aware QG or answer-agnostic QG [130]) and generating a question based on the statement.
Datasets. QG and QA are dual tasks [131]. They both require reasoning between questions and answers. As a result, some MCQ datasets initially designed for QA tasks, including SciQ [132], RACE [133], and FairytaleQA [134], are also leveraged for researches on QG [135, 136, 137, 138]. Besides, there are some datasets particularly constructed for QG, such as LearningQ [139], KHANQ [140], and EduQG [141]. These datasets extend across various subjects (e.g., science, medicine) and various education levels (e.g., preliminary school, middle school, and university.), the details of which are displayed in Table II. Regarding these datasets, LearningQ is collected from instructor-crafted learning questions, which contain cognitively demanding documents for QG that requires reasoning, but the absence of answers limits its application to answer-aware QG. KHANQ provides (context, prompt, question) triples collected from questions about the online courses raised by students. The prompt, which indicates the background knowledge of students or presupposition of the questions, provides hints on the information that should be brought into focus during question generation. EduQG is a high-quality MCQ dataset generated by educational experts and questions are linked to their cognitive complexity, which makes this dataset move closer to reality.
Dataset Science subject Source Question type Answer type #Q
SciQ science online science normal MC& 13.713.713.713.7K
textbooks extractive
RACE linguistic secondary school normal& MC& 100100100100K
English exams cloze extractive
FairytaleQA literature grade K-8 normal extractive 10101010K
narrative stories
LearningQ computing/ online education normal - 231231231231K
science/business/ platform
@@ 3.2.1 Question Generation
KHANQ science online education normal extractive 1111K
platform
EduQG politics/physiology/ online textbooks normal& MC& 3333K
biology/business/ cloze extractive
@@
MCQL biology/physics/ website normal MC 7.17.17.17.1K
chemistry
Televic math/health/ online education normal MC 62626262K
history/geography/ platform
@@
TABLE II: Comparison of QG and DG datasets. The format of the question is categorized as either normal (e.g., interrogative sentence.) or cloze (e.g., fill-in-the-blank.), and the answer is categorized as either multiple-choice (MC) or extractive (parts of the text are considered as the answer) [141].
@@
Methods. Early methods for general QG primarily relied on rule matching [49]. Du et al. [50] first applied the Seq2Seq model with attention to automatic QG. Built upon this framework, methods such as integrating linguistic features [51, 52, 53], training a multi-modal QG model [54], leveraging multi-task learning [55] and reinforcement learning [56] have also been augmented to optimize the QG models. With the advent of language models, the BERT model and LLM GPT-3 have been applied to QG [142, 143]. Besides, regarding answer information in the QG system, researchers proposed to encode the answer together with the context [144], utilize the position information of the answer in the context [145, 146], and achieve answer-aware QG, while using summaries of the texts as input can serve as a strategy for answer-agnostic QG as it can help the modeling of topic salience, which denotes the importance of specific words or phrases in context [130].
Importantly, there are some unique notes when we adopt QG to the education domain. First, controlling the difficulty levels of the generated questions is vital for improving the efficiency of education. Gao et al. [57] proposed a method for difficulty-controllable QG, and the difficulty of a question is judged according to whether a well-trained QA model is able to answer it correctly. Then Cheng et al. [58] redefined the difficulty level as the number of inference steps required for a question to answer. Recently, Uto et al. [59] considered the difficulty associated with learners’ ability, which is measured by the item response theory [147]. Second, it is important to align generated questions to the syllabus, which benefits to identify the focus of the test quiz. Hence, multiple studies [60, 61] train binary classifiers or ranking models to judge whether the generated question is related to certain entities and concepts or pedagogical contents with the objective of maximizing both relevance and diversity. Third, customized education is in high demand and we notice there are a few studies focusing on generating personalized questions for students. Srivastava et al. [62] developed a knowledge tracking model that uses students’ answer histories to predict their performance on new questions, fine-tuning it with an autoregressive language model. Subsequently, Wang et al. [63] introduced the few-shot knowledge tracking model to predict question difficulty by incorporating sequences of student states and questions.

### v4

@@ 3.1 Question Answering
Dataset Science subject Source Context #Q Top 1 Top 2 Top 3
[…]
physical science science curricula Llama-2 RAG [85] SSCGN [86] ISAAQ [87]
[…]
PMI [89] ESIM [90] IR [89]
[…]
Claude 3.5 Sonnet GPT-4o Gemini-1.5-Pro
[…]
language science science curriculum text Multimodal-T-SciQ[18] MC-CoT-F[92] Honeybee[93]
[…]
board exam MedPaLM-2 GPT-4 Flan-PaLM
[…]
VOD BioLinkBert Med-Palm Flan-PaLM
[…]
EE/CS GPT-4 LLaMA3-70B Qwen1.5-110B
[…]
DMR [97] SMR [97] MaGNET [98]
[…]
EPT-XL [99] GEO [100] EPT-L [99]
[…]
Teaching-GPT-4 [101] Multi-view [102] Generate&Rank [103]
[…]
ELASTIC [104] Exp-Tree [19] Multi-view [102]
[…]
Self-Refine [105] Toolformer [106] GPT-3
[…]
DUP-GPT-4 [107] CoT-GPT-4 [27] PS-GPT-4 [108]
[…]
InstructBLIP [109] LLaMA-2-Chat-7B Vicuna-13B
[…]
Table-LLaVA 13B [110] Vicuna-1.5 7B Llama2+Oracle
[…]
Even though QA has been explored as a popular NLP task for decades, educational QA requires a system to answer the domain-specific questions, related to natural science, social science, language science and so on [111]. These systems act as assistants to help answer difficult questions such that the students could learn from them and boost their performance. Question answering in different subjects may be presented in different formats. In Figure 4, we display two typical examples of question answering in the education domain. To answer the displayed questions, the systems need domain-specific knowledge like mathematics and science with the given contexts like text and images. The output of the questions can be presented as an entity or a math expression. In this section, we will present cutting-edge techniques for Textbook QA and MWP solving.
@@ 3.1.1 Textbook Question Answering
FIGURE 4: The examples of question answering in education domain. The red box denotes the inputs and the green box denotes the required outputs for the tasks. The displayed examples are extracted from SCIENCEQA and GSM8K, respectively.
@@
Textbook Question Answering (TQA) is a task that requires a system to comprehensively understand the multi-modal information from the textbook curriculum, spreading across text documents, images, and diagrams. We formulate the general loss function of TQA as follows for simplicity:
$L_{\text{TQA}}=-\mathbb{E}_{(P,Q,A)\in\mathcal{D}}\log P_{\theta}(A|Q,P)$
where $P,Q,A$ are contexts, questions and answers from data collection $\mathcal{D}$, respectively. $P$ may be presented in non-textual modalities. The major challenge of TQA is to comprehend the multi-modal domain-specific contexts as well as the questions, and then identify the key information to the questions.
@@
Datasets. Kembhavi et al. [112] introduced TQA dataset, which aims to evaluate a system that contains multi-modal contexts and diverse topics in the scientific domain. Similar datasets such as AI2D [113], SCIENCEQA [1] are other large-scale TQA datasets with annotated lectures and explanations covering diverse science topics including natural science, social science, and language science. Some other datasets such as GeoSQA [88], MedQA [94], MedMCQA [114] cover specific subjects such as geography and medicine. A recently released TheoremQA dataset further covers textbook questions at the university level [2].
Methods. From the technical perspective, TQA can be deemed as Visual Question Answering (VQA) in nature [115, 116, 117]. Conventional VQA studies employ deep-learning models to encode the questions and images [118, 119]. Then the multi-modal information is fused together to comprehend the questions. While there are significant similarities between VQA and TQA, we highlight some specific obstacles and limitations involved due to the distinct educational scenarios.
  • The subject-specific QA data is limited for training. Even though PLMs are frequently utilized as the backbone model, they are short of the subject knowledge. To bridge the gap, some studies [11, 12] fine-tune the contextual knowledge in the dataset as language modeling, which boosts the performance of TQA when only a small number of examples are available in the specific domains:
    $L_{\text{PreTrain}}=-\mathbb{E}_{P\in\mathcal{D}}\log P_{\theta}(P_{m}|Q,P_{\backslash m}),$
[…]
    $L_{\text{TQA}}=-\mathbb{E}_{(P,Q,A)\in\mathcal{D}\bigcup\tilde{\mathcal{D}}}\log P_{\theta}(A|Q,P),$
    where $\tilde{\mathcal{D}}$ is augmented data set from different subjects. $P_{\theta}$ is a VQA model trained with adversarial training to learn subject-invariant features [13].
  • The contexts may be presented in diverse formats. To understand the diagrams and tables, graph-based parsing methods are developed to extract the concepts from diagrams [113]. Optical Character Recognition (OCR) is employed to identify the answers from the charts, which is further aligned with the questions [14, 15]. These methods conduct pre-processing:
    $\tilde{P}=\mathcal{T}(P).$
    Here, $\mathcal{T}$ is a set of off-the-shelf toolkits for calling. After that, $\tilde{P}$ will be deemed as the contexts of $Q$.
  • Zero-shot issues frequently occur. TQA in education has a higher requirement for reasoning under zero-shot scenarios. Recently, LLMs have shown impressive zero-shot capabilities in various NLP tasks, a number of studies [1, 18, 16] applied GPT-3 to solve the TQA. To apply the full advantage of LLMs to image modality, they use caption models to translate visual information into language modality. The paradigm of TQA systems becomes:
[…]
    Liu et al. [17] extended LLMs to multi-modalities, which results in a multi-modal model that connects a vision encoder and an LLM for general-purpose visual and language understanding:
    $A=\text{MLLM}(Q,P).$
    More recent studies [32, 2] develop compositional reasoning with LLMs as a planner that assembles a sequence of tools (e.g., Hugging Face as image captioner, Github as text detector, Web Search as bing search, and Python as program verifier.) to execute and answer the textbook questions. We can formulate their methods as:
    $A=\text{LLM}(Q,P,\mathcal{T}).$
The above methods could effectively solve part of the challenges, but based on Table II, there is still room for improvement and we observe a severe performance gap of TQA tasks between the mainstream and the non-mainstream subjects such as geography and medicine, which is caused by the imbalanced collection of QA data. This could be solved by constructing larger and wider scale of educational data.
@@ 3.1.2 Math Word Problem solving
Math Word Problem (MWP) solving is a typical category of QA specifically in mathematics subjects, aiming to convert a narrative description to an abstract expression:
$L_{\text{MWP}}=-\mathbb{E}_{(Q,E)\in\mathcal{D}}\log P_{\theta}(E|Q).$
@@
Here, $Q$ and $E$ are a question and its corresponding math expression, which consists of a sequence of tokens and symbols, respectively. This task is challenging because there remains a wide semantic gap in parsing human-readable words into machine-understandable logic.
@@
Datasets. Dolphin-1818[120] is an early large-scale MWP solving dataset, which contains over $18,000$ annotated math word problems of elementary mathematics. DRAW-11[121] is introduced as a testbed for algebra word problem solvers. Math2323[122] is a Chinese MWP solving dataset. Subsequently, MathQA [123] and ASDiv [124] further extend the scale and diversity of the existing datasets, which makes it particularly suited for usage in training deep learning models to solve MWP problems. Cobbe et al. [125] introduced a more challenging dataset GSM88K, which is complicated in high linguistic diversity. IconQA [126] extends MWP solving to abstract diagram understanding and comprehensive cognitive reasoning by involving diagrams in problems. Moreover, TABMWP [127] includes tables to enhance the difficulty of the MWPs, which requires the table parsing and understanding capabilities [127]. As we can see, there is a set of MWP datasets that examine various reasoning capabilities. With the development of this task, the evaluated datasets become increasingly complicated in linguistic diversity and multi-modalities.
@@
Methods. MWP solving is a popular NLP application in the education domain for granted since it can be utilized as a teaching assistant in math. The investigation of MWP solving can be traced back to the 19601960[128]. Early work leverages manually crafted rules and schemas for pattern matching, where human intervention is heavily relied on. Subsequently, various strategies of feature engineering and statistical learning are proposed to boost the performance. With the development of techniques and the emergence of large-scale datasets, deep learning-based methods have been widely applied to MWP solving. The majority of the methods follow the paradigm of machine translation [122], which first encodes the narratives, then decodes the mathematical expression token as the output sequence. Compared with machine translation, there are some distinct solutions that have been proposed due to the structured outputs of MWPs.
  • Some MWPs are complex in logic. To improve the generation of complex solution expressions, some methods are proposed to formulate a mathematical expression as an abstract tree [19, 20, 21, 22]. To improve the understanding of the narratives, some studies focus on improving the encoding of the narratives by re-organizing the textual contexts with structure [23, 24], which may increase computation complexity as a drawback. In this case, graph networks are frequently involved in:
    $\mathcal{E}=P_{\theta}(E|Q)$ $=\text{GNN}(\mathcal{Q}).$
    Here, equations and questions can be represented as dependency trees $\mathcal{Q}$ and abstract syntax trees $\mathcal{E}$, respectively.
  • There is a lack of annotated expressions in the MWPs. Due to the weak supervision caused by the lack of annotated expressions, multiple studies [19, 25] model the MWP solving task as follows:
    $L_{\text{MWP}}=-\mathbb{E}_{(Q,A)\in\mathcal{D}}\log\sum_{E}P(A|E)P_{\theta}(E|Q).$
    Also, some researchers [26] apply knowledge distillation to MWP solving tasks, which learns a smaller model from a large pre-trained generic model but may also include some unexpected noise.
@@
  • Zero/few-shot complex MWPs in real-world scenarios. MWP solving usually encounters problems with complex reasoning steps under zero-shot. Chain-of-Thought (CoT) is a typical prompting strategy that is meant to decompose a complicated MWP into sub-questions such that LLMs could conduct reasoning step by step [27]. Follow-up methods improve it by designing more interactive prompting or constructing more efficient demonstrations [28, 29, 30, 31, 32]. Similar to the TQA tasks, researchers [33] leverage LLMs as an agent [129] and delegate the computation to a program interpreter, which can be conducted by the LLMs and executed by a Python interpreter.
We observe the performance of MWP solving has a large variance across different datasets in Table II. With the increasing complexity of the modalities and questions, the tasks become more challenging. Hence, more advanced and powerful methods should be explored to improve the performance.
@@ 3.2 Question Construction
Dataset Science subject Source Q&A type #Q Top 1 Top 2 Top 3
[…]
textbooks MC&extractive EduQG [130] Leaf [132] DDS+ST [131]
[…]
English exams MC&extractive BDG [133] GPT DDS+ST [131]
[…]
science/business/ platform - Attention [135] Seq2Seq [136] H&S [137]
@@
[…]
platform extractive T5 [139] BART [140] UniLM [141]
[…]
economics/science/ textbooks extractive T5 [139] BART [140] MBART [143]
environmental studies
[…]
biology/business/ MC&extractive DCS [41] T5 [139]
@@
[…]
chemistry MC LR+RF [145] LR+LM [145] RF [145]
[…]
history/geography/ platform MC ChatGPT DQ-SIM [146] mT5 [147]
@@
[…]
3.2.1 Question Generation
The goal of QG is to automatically generate a question from a given sentence or paragraph. The challenge lies in identifying the key statement based on the context with or without the utilization of the answer (namely answer-aware QG or answer-agnostic QG [149]) and generating a question based on the statement, which can be described as follows:
$L_{\text{QG}}=-\mathbb{E}_{(C,[A],Q)\in\mathcal{D}}\log P_{\theta}(Q\mid C,[A])$
Here $C$ and $Q$ are context and the generated question, respectively. $[A]$ is the answer which is optional to the input.
Datasets. QG and QA are dual tasks [150]. They both require reasoning between questions and answers. As a result, some MCQ datasets initially designed for QA tasks, including SciQ [151] and RACE [152], are also leveraged for research on QG [153, 154, 155, 156]. Besides, there are some datasets particularly constructed for QG, such as LearningQ [134], KHANQ [138], EduQG [144] and EduProbe [142]. These datasets extend across various subjects (e.g., science, medicine) and various education levels (e.g., preliminary school, middle school, and university.), the details of which are displayed in Table II. Regarding these datasets, LearningQ is collected from instructor-crafted learning questions, which contain cognitively demanding documents for QG that require reasoning, but the absence of answers limits its application to answer-aware QG. KHANQ and EduProbe provide (context, textual cue, question) triples. EduQG is a high-quality MCQ dataset generated by educational experts and questions are linked to their cognitive complexity, which makes this dataset move closer to reality.
@@ 3.2.1 Question Generation
Methods. Early methods for general QG primarily relied on rule matching [157]. Du et al. [158] first applied the Seq2Seq model with attention to automatic QG. Built upon this framework, methods such as integrating linguistic features [159], training a multi-modal QG model [160], leveraging multi-task learning [161] and reinforcement learning [162] have also been used to optimize the QG models. With the advent of language models, the BERT model and GPT-3 have been applied to QG [163, 164]. Besides, regarding answer information in the QG system, researchers proposed to encode the answer together with the context [165], utilize the position information of the answer in the context [166], and achieve answer-aware QG, while using summaries of the texts as input can serve as a strategy for answer-agnostic QG as it can help the modeling of topic salience, which denotes the importance of specific words or phrases in the context [149]. Importantly, there are some unique challenges when we adopt QG systems to the education domain.
  • The generated questions ignore the difficulty stratification. Some general methods lack an effective way to control the difficulty of the question, which is vital for improving the efficiency of education. The difficulty-controllable QG task can be described as:
    $L_{\text{QG}}=-\mathbb{E}_{(C,A,Q,D)\in\mathcal{D}}\log P_{\theta}(Q\mid C,A,D)$
    Here $D$ represents the difficulty levels, which can be formulated in several ways: whether a well-trained QA model can answer it [34], the number of inference steps [35], and the educational paradigms like Item Response Theory [36] or Bloom’s Taxonomy [37] that categorize learning objectives into cognitive, affective, and psychomotor domains. Recently, Hwang et al. [167] applied few-shot prompting of GPT-4 to generate questions aligned with Bloom’s Taxonomy [37]. Lee et al. [38] and Maity et al. [39] also designed an effect prompt for the QG system to cover a wide range of difficulty levels.
  • The generated questions show weak relevance to the syllabus. Aligning generated questions to the syllabus can be beneficial for identifying the focus of the test quiz, which can be achieved by including a selection step for content in general QG methods. Multiple studies [40, 41] train binary classifiers or ranking models to judge whether the generated question is related to concepts or pedagogical contents with the objective of maximizing both relevance and diversity.
    $P_{\phi}(K\mid Q,C)=\text{Classifier}(Q,C,K),$
    where $K$ represents certain keywords or topics of the syllabus. Besides, Xiao et al. [42] generated questions that align with given topic keywords using the plug-and-play language models [168].
  • The generated questions should be customized to students. QG systems for customized education focus on generating personalized questions for students. Srivastava et al. [43] developed a knowledge-tracking model that uses students’ answer histories to predict their performance on new questions, fine-tuning it with an autoregressive language model. Subsequently, Wang et al. [44] introduced the few-shot knowledge tracking model to predict question difficulty by incorporating sequences of student states and questions.
    $L_{\text{QG}}=-\mathbb{E}_{(C,S,Q,D)\in\mathcal{D}}\log P_{\theta}(Q\mid C,S,D)$
[…]
Even though various methods have been proposed, question generation is still a difficult task to ensure its alignment to the real world. More studies focusing on developing syllabus-guided and customized questions are highly-anticipated.